40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 1 08:50:48 2022, yuta, Summary, LSC, FPMI with REFL/AS55 trial 
    Reply  Thu Jul 14 00:10:08 2022, yuta, Summary, LSC, FPMI with REFL/AS55 trial continued AS55_I.pngAS55_Q.pngREFL55_I.pngREFL55_Q.png
       Reply  Fri Jul 15 19:13:22 2022, Paco, Summary, LSC, FPMI with REFL/AS55 demod phase adjust 
          Reply  Fri Jul 15 22:36:04 2022, rana, Summary, LSC, FPMI with REFL/AS55 demod phase adjust 
          Reply  Mon Jul 18 16:39:07 2022, Paco, Summary, LSC, FPMI locking procedure using REFL55 and AS55 DARM_07_18_2022_FMPI.pdfCARM_07_18_2022_FPMI.pdfMICH_07_18_2022_FPMI.pdffpmi_darm_nb_2022_07.pdf
             Reply  Mon Jul 18 21:41:42 2022, Anchal, Summary, LSC, FPMI locking procedure using REFL55 and AS55 
             Reply  Tue Aug 9 19:54:31 2022, yuta, Summary, LSC, FPMI locking tonight 
                Reply  Thu Aug 18 14:49:35 2022, Yehonathan, Summary, LSC, FPMI Sensitivity Sensitivity_Plot_1344133503.pdf
                   Reply  Thu Aug 18 18:10:49 2022, Koji, Summary, LSC, FPMI Sensitivity 
Message ID: 17012     Entry time: Mon Jul 18 16:39:07 2022     In reply to: 17007     Reply to this: 17016   17069
Author: Paco 
Type: Summary 
Category: LSC 
Subject: FPMI locking procedure using REFL55 and AS55 

[Yuta, Paco]

In summary, we locked FPMI using REFL55_I, REFL55_Q, and AS55_Q. The key to success was to mix POX11_I and POY11_I in the right way to emulate CARM/DARM, and to find out the correct demodulation phase for AS55.


Procedure

  1. Close PSL shutter and zero offsets in AS55, REFL55, POX11, POY11, and ASDC
    • For ASDC run python3 resetOffsets.py -c C1:LSC-ASDC_IN1, otherwise use the zer offsets on I and Q inputs from the RFPD medm screen.
  2. Lock XARM/YARM using POX/POY to tune demodulation phase.
    • Today, the demode phase in POX11 changed to 104.801, and POY11 to -11.256 deg.
  3. XARM and YARM are used in the following configuration
    • INMAT
      • 0.5 * POX11_I - 0.5 * POY --> XARM
      • 0.5 * POX + 0.5*POY --> YARM
      • REFL55_Q --> MICH (** this should be turned on after POX11/POY11)
    • LSC Filter gains
      • XARM = 0.012
      • YARM = 0.012
      • MICH = +40 (note the sign flip from last time)
    • OUTMAT
      • XARM --> 0.5 * ETMX - 0.5 * ETMY
      • YARM --> MC2
      • MICH --> BS
    • UGFs (sanity check)
      • XARM (DARM) ~ 100 Hz
      • YARM (CARM) ~ 200 Hz
      • MICH (MICH) ~ 40 Hz
  4. Run MICHOpticalGainCalibration.ipynb to see if ASDC vs REFL55_Q looks nice (ellipse in the XY plot), and find any residual offset in REFL55_Q.
    • If the plot doesn't look nice in this regard, the IFO needs to be aligned.
  5. Sensing matrix for CARM/DARM and MICH.
    • With the DARM, CARM and MICH lines on, verify the demod error signals look ok both in mag and phase.
    • For example, we found that CARM error signals were correctly represented by either 0.5 * POX11_I + 0.5 * POY11_I or 0.5 * REFL55_I.
    • Similarly, we found that DARM error signal was correctly represented by either 0.5 * POX11_I - 0.5 * POY11_I or 2.5 * AS55_Q.
    • To find this, we minimized CARM content in AS55_Q, as well as CARM content in REFL55_Q.
  6. We acquired the lock by re-configuring the error point as below:
    • INMAT
      • 0.5*REFL55_I --> YARM (CARM)
      • 2.5 * AS55_Q --> XARM (DARM)
    • During the hand-off trials, we repeatedly ran the sensing matrix and UGF measurements while stopping at various intermediate mixed error points to check how the error signal calibrations changed if at all.
      • Attachment #1 shows the DARM OLTF using POX/POY (blue), only with CARM handoff (green), and after DARM handoff (red)
      • Attachment #2 shows the CARM OLTF using POX/POY (blue), only with CARM handoff (green), and after DARM handoff (red)
      • Attachment #3 shows the MICH OLTF using POX/POY (blue), only with CARM handoff (green), and after DARM handoff (red)
    • The sensing matrix after handoff is below:
Sensing Matrix with the following demodulation phases
{'AS55': 192.8, 'REFL55': 95.63177865911078, 'POX11': 104.80089727128349, 'POY11': -11.256509422276006}
Sensors          	           DARM     	           CARM     	            MICH     	
C1:LSC-AS55_I_ERR_DQ	5.09e-02 (89.6761 deg)	2.03e-01 (-114.513 deg)	1.28e-04 (-28.9254 deg)	
C1:LSC-AS55_Q_ERR_DQ	4.78e-02 (88.7876 deg)	3.61e-03 (-68.7198 deg)	8.34e-05 (-39.193 deg)	
C1:LSC-REFL55_I_ERR_DQ	5.18e-02 (-92.2555 deg)	1.20e+00 (65.2507 deg)	1.15e-04 (-102.027 deg)	
C1:LSC-REFL55_Q_ERR_DQ	1.81e-04 (59.0854 deg)	1.09e-02 (-114.716 deg)	1.77e-05 (-23.6485 deg)	
C1:LSC-POX11_I_ERR_DQ	8.51e-02 (91.2844 deg)	4.77e-01 (67.1709 deg)	7.97e-05 (-72.5252 deg)	
C1:LSC-POX11_Q_ERR_DQ	2.63e-04 (114.584 deg)	1.32e-03 (-113.505 deg)	2.10e-06 (118.146 deg)	
C1:LSC-POY11_I_ERR_DQ	1.58e-01 (-88.9295 deg)	6.16e-01 (67.6098 deg)	8.71e-05 (172.73 deg)	
C1:LSC-POY11_Q_ERR_DQ	2.89e-04 (-89.1114 deg)	1.09e-03 (70.2784 deg)	3.77e-07 (110.206 deg)	

Lock gpstimes:

  1. [1342220242, 1342220260]
  2. [1342220420, 1342220890]
  3. [1342221426, 1342221574]
  4. [1342222753, 1342223230]

Sensitivity estimate (NANB)

Using diaggui, we look at the AS55_Q error point and the DARM control point (C1:LSC-XARM_OUT). We roughly calibrate the error point using the sensing matrix element and actuation gain at the DARM oscillator freq 4.78e-2 / (10.91e-9 / 307.880^2). The control point is calibrated with a 0.95 Hz SUS pole. Attachment #4 shows the sensitivity estimate.

Attachment 1: DARM_07_18_2022_FMPI.pdf  17 kB  Uploaded Mon Jul 18 18:44:54 2022  | Hide | Hide all
DARM_07_18_2022_FMPI.pdf DARM_07_18_2022_FMPI.pdf DARM_07_18_2022_FMPI.pdf
Attachment 2: CARM_07_18_2022_FPMI.pdf  17 kB  Uploaded Mon Jul 18 18:45:01 2022  | Hide | Hide all
CARM_07_18_2022_FPMI.pdf CARM_07_18_2022_FPMI.pdf CARM_07_18_2022_FPMI.pdf
Attachment 3: MICH_07_18_2022_FPMI.pdf  27 kB  Uploaded Mon Jul 18 18:45:10 2022  | Hide | Hide all
MICH_07_18_2022_FPMI.pdf MICH_07_18_2022_FPMI.pdf MICH_07_18_2022_FPMI.pdf
Attachment 4: fpmi_darm_nb_2022_07.pdf  46 kB  Uploaded Mon Jul 18 18:45:23 2022  | Hide | Hide all
fpmi_darm_nb_2022_07.pdf
ELOG V3.1.3-